10x-2y+z=12,x+9y-z=10,2x-y+11z=20 i need answer fr tz question in gauss elimination method and gauss jourdan method imediately ..its so urgent
Answer by Edwin McCravy(20066)

The correct answer, by using a computer program, is 

x = 178/141
y = 1144/987
z = 1672/987

No respectable teacher would give you a problem
to do by Gaussian or Gauss-Jordan method or ANY
method BY HAND that has such a god-awful solution 
as that. You're bound to have made a typo. If you'll 
fix the typo you made, we'll help you.  We're not
going to manipulate all the terrible fractions that 
it would take to do that by hand. It would take an 
hour or more.

Edwin